OpenStudy (anonymous):

Write the equation 2x - 3y = 6 in slope-intercept form. A.) y =x + 3 B.) y =x + 3 C.) y =x - 2

OpenStudy (anonymous):

B

OpenStudy (johnweldon1993):

So isolate 'y' on the left side...how would you do that?

OpenStudy (anonymous):

By subtraction the 3 from both sides.

OpenStudy (anonymous):

* Subtracting

OpenStudy (anonymous):

I'm sorry im missing 2/3x in all the problems

OpenStudy (johnweldon1993):

Not quite....you would subtract the 2x from both sides 2x - 3y = 6 -2x -2x -3y = 6 - 2x Then you would divide everything by -3 to isolate 'y' -3y = 6 - 2x ------------ -3 -3 -3 y = -2 + 2/3x or written correctly y = 2/3x - 2

OpenStudy (anonymous):

Hmmmmm Good work

OpenStudy (johnweldon1993):

see how that was done @DvonCruse ?

OpenStudy (anonymous):

I see now. Thanks

OpenStudy (anonymous):

@johnweldon1993